OpenStudy (anonymous):

What is the quadratic function that is created with roots at 2 and 4 and a vertex at (3, 1)? please show work

OpenStudy (perl):

your quadratic equation is y = a ( x - 2)( x - 4) and you are given vertex , you can plug this in to solve for 'a'

OpenStudy (perl):

1 = a ( 3 -2 ) ( 3 - 4)

OpenStudy (anonymous):

okay the roots are given so the quadratic equation can be written as y=a(x-2)(x-4) y=a(x^2-4x-2x+8)=a(x^2-6x+8) now we have to find value of constant a vertex=(3,1) I pu x=3 and y=1 in the above equations and get the value of a so I got 1=a(3^2-6 x 3 +8) a=-1 so the quadratic equation is y=-(x^2-6x+8) is this correct?^^

OpenStudy (perl):

yes i got a = -1 as well

OpenStudy (anonymous):

so y=-(x^2-6x+8) is my final answer, correct?

OpenStudy (perl):

you should distribute the negative sign, might as well

OpenStudy (anonymous):

can you send me the final answer after distributing?

OpenStudy (perl):

y = -x^2 + 6x - 8

OpenStudy (anonymous):

thank u:)